### v1

@@ 4 Characterizing the Natural Filter
Fix a family of non-degenerate PLDs $\mathcal{L}$, i.e. every $L\in\mathcal{L}$ is supported at at least two points in $\mathbb{R}$. The natural filter Algorithm˜1 is free for every budget $B$ and every query capacity $k$ if and only if $\mathcal{L}^{\infty}$ is well-ordered with respect to $\preceq$.
We prove here that families of queries that are well-ordered under composition satisfy the commutativity condition. The proof of the other direction is significantly more involved, so we delay it until Section˜6.
Proof of Sufficient Condition for Theorem˜3.
Assume that $\mathcal{L}^{\infty}$ is well-ordered and let $L\in\mathcal{L},\mathcal{L}^{\prime}\subseteq\mathcal{L}^{<k}\subseteq\mathcal{L}^{\infty}$. For convenience, we will denote by $h_{L}$ the unique hockey-stick curve associated to a PLD $L$, namely $h_{L}(x)=\mathop{\mathbb{E}}_{Z\sim L}[(1-xe^{-Z})_{+}]$ (see Proposition˜5). Now, as in the proof of Theorem˜2, we already have $\sup\{L\oplus\mathcal{L}^{\prime}\}\preceq L\oplus\sup{\mathcal{L}^{\prime}}$ for free, so due to Proposition˜11, we just need to show that $h_{L\oplus\sup{\mathcal{L}^{\prime}}}\preceq h_{\sup\{L\oplus\mathcal{L}^{\prime}\}}$. To that end, let $\gamma>0$. By well-ordering of $\mathcal{L}^{\prime}$ and by Proposition˜11, we can find some $L^{\prime}\in\mathcal{L}^{\prime}$ so that $h_{\sup{\mathcal{L}^{\prime}}}\preceq h_{L^{\prime}}+\gamma$. In this case, we have
$h_{L\oplus\sup{\mathcal{L}^{\prime}}}(x)$ $=\mathop{\mathbb{E}}_{Z\sim L}[\mathop{\mathbb{E}}_{Z^{\prime}\sim\sup{\mathcal{L}^{\prime}}}[(1-xe^{-(Z+Z^{\prime})})_{+}]]$
$=\mathop{\mathbb{E}}_{Z\sim L}[h_{\sup{\mathcal{L}^{\prime}}}(xe^{-Z})]$
$\leq\mathop{\mathbb{E}}_{Z\sim L}[h_{L^{\prime}}(xe^{-Z})+\gamma]$
$=\mathop{\mathbb{E}}_{Z\sim L}[\mathop{\mathbb{E}}_{Z^{\prime}\sim L^{\prime}}[(1-xe^{-(Z+Z^{\prime})})_{+}]]+\gamma$
$=h_{L\oplus L^{\prime}}(x)+\gamma$
$\leq h_{\sup\{L\oplus L^{\prime}\}}(x)+\gamma$
for any $x\in\mathbb{R}^{\times}$, which yields the desired result as $\gamma$ was arbitrary. ∎
@@ 5 Notable implications
5.1 PLD and $f$-DP filters are NOT free in general
Theorems˜2 and 3 give a necessary and sufficient condition for free PLD filters. We now switch our view to $f$-DP using Propositions˜10 and 11, for which free filters correspond to Algorithm˜2 in Appendix, to illustrate these results and their implications. The condition in Theorem˜2, which is necessary and sufficient for free privacy filters, translates as:
$\forall f\in\mathcal{F},\ \text{and}\ \mathcal{F}^{\prime}\subseteq\mathcal{F}^{\infty},f\otimes\operatorname{conv}(\mathcal{F}^{\prime})=\operatorname{conv}(f\otimes\mathcal{F}^{\prime}),$ (1)
where $\otimes$ is $f$-DP composition, $f\otimes\mathcal{F}^{\prime}=\{f\otimes f^{\prime}:f^{\prime}\in\mathcal{F}^{\prime}\}$, and $\operatorname{conv}$ is the lower convex envelope of a set of tradeoff curves.
All examples also apply to PLD by Proposition˜10, but we visualize them using trade-off curves. For examples and visualizations, we work with piece-wise linear tradeoff curves, which can be composed exactly using the following result.
@@ 5.1 PLD and $f$-DP filters are NOT free in general
Consider two piecewise-linear tradeoff curves $f_{1},f_{2}$. Denote $\alpha_{i,n_{i}}\in[0,1]$ the lengths of intervals over which $f_{i}$ is linear, and $f^{\prime}_{i,n_{i}}\leq 0$ the slope of $f_{i}$ on those intervals. Denote $\delta_{i}\triangleq 1-f_{i}(0)$.
That is, $\sum_{j=1}^{n_{i}}\alpha_{i,j}=1$, and, noting $m_{i}(\alpha)=\arg\max_{n:\in\{1,n_{i}\}}\sum_{j=1}^{n}\alpha_{i,j}\leq\alpha$, we have:
$f_{i}(\alpha)=1-\delta_{i}+\sum_{j=1}^{m_{i}(\alpha)}f^{\prime}_{i,j}\alpha_{i,j}+\big(\alpha-\sum_{j=1}^{m_{i}(\alpha)}\alpha_{i,j}\big)f^{\prime}_{i,j+1}.$
Let $k\in[1,n_{1}n_{2}]$ index all combinations of $(j_{1,k},j_{2,k})\in[1,n_{1}]\times[1,n_{2}]$ in order of increasing values of $f^{\prime}_{1,j_{1,k}}f^{\prime}_{2,j_{2,k}}$. That is, $k$ indexes all products of a slope from $f_{1}$ and a slope for $f_{2}$ (which are now positive) sorted in decreasing order of magnitude. With ne minus sign, the first index is thus the steepest negative slope.
Then the composition $f\triangleq f_{1}\otimes f_{2}$ is also a piecewise linear tradeoff curve, with $\delta=\delta_{1}+\delta_{2}-\delta_{1}\delta_{2}$, and $\alpha_{k}=\alpha_{1,k}\alpha_{2,k}$ and slopes $f^{\prime}_{k}=-f^{\prime}_{1,j_{1,k}}f^{\prime}_{2,j_{2,k}}$. That is, denoting $m(\alpha)=\arg\max_{k:\in\{1,n_{1}n_{2}\}}\sum_{k=1}^{n}\alpha_{k}\leq\alpha$, $f=f_{1}\otimes f_{2}$ is such that:
$f(\alpha)=1-\delta_{1}-\delta_{2}+\delta_{1}\delta_{2}+\sum_{j=1}^{m(\alpha)}f^{\prime}_{k}\alpha_{k}+\big(\alpha-\sum_{j=1}^{m(\alpha)}\alpha_{k}\big)f^{\prime}_{k+1}.$

### v2

@@ 5 Characterizing the Natural Filter
Let $\mathcal{L}\subseteq\mathcal{L}_{\mathrm{all}}$ and $k\in\mathbb{Z}^{+}\cup\{\infty\}$. If $\mathcal{L}^{<k}$ is totally-ordered, then $\operatorname{PLD}(\operatorname{Nat}_{\mathcal{L},B}^{k})\preceq B$ for every $B\in\mathcal{L}_{\mathrm{all}}$. Conversely, if the non-degenerate members of $\mathcal{L}^{<k}$ are not totally-ordered, then $\operatorname{PLD}(\operatorname{Nat}_{\mathcal{L},B}^{2k-3})\npreceq B$ for some choice of $B\in\mathcal{L}_{\mathrm{all}}$.
Proving the forward direction is easy. We show in general that the supremum commutes with convolution for totally-ordered families of queries.
Lemma 3.
Let $L$ be a PLD and let $\mathcal{L}$ be a totally-ordered set of PLDs. Then $L\oplus\sup{\mathcal{L}}=\sup(L\oplus\mathcal{L})$.
Proof.
As in the proof of Theorem˜7, we already have $\sup(L\oplus\mathcal{L})\preceq L\oplus\sup{\mathcal{L}}$ for free, so due to Proposition˜11, we just need to show that $h_{L\oplus\sup{\mathcal{L}}}\preceq h_{\sup(L\oplus\mathcal{L})}$. By total-ordering of $\mathcal{L}$ and by Proposition˜11, we can find a sequence $L_{1}\preceq L_{2}\preceq\dots\in\mathcal{L}$ so that $h_{L_{t}}\to h_{\sup{\mathcal{L}}}$ pointwise. For any $x\in\mathbb{R}^{\times}$, monotone convergence yields
$h_{L\oplus\sup{\mathcal{L}}}(x)$ $=\mathop{\mathbb{E}}_{Z\sim L}[\mathop{\mathbb{E}}_{Z^{\prime}\sim\sup{\mathcal{L}}}[(1-xe^{-(Z+Z^{\prime})})_{+}]]$
$=\mathop{\mathbb{E}}_{Z\sim L}[h_{\sup{\mathcal{L}}}(xe^{-Z})]$
$=\mathop{\mathbb{E}}_{Z\sim L}[\lim_{t\to\infty}h_{L_{t}}(xe^{-Z})]$
$=\lim_{t\to\infty}\mathop{\mathbb{E}}_{Z\sim L}[h_{L_{t}}(xe^{-Z})]$
$=\lim_{t\to\infty}\mathop{\mathbb{E}}_{Z\sim L}[\mathop{\mathbb{E}}_{Z^{\prime}\sim L_{t}}[(1-xe^{-(Z+Z^{\prime})})_{+}]]$
$=\lim_{t\to\infty}h_{L\oplus L_{t}}(x)$
$\leq h_{\sup(L\oplus\mathcal{L})}(x).$
Proving the other direction is substantially more challenging. The key lemma is that, given a pair of non-degenerate queries that are not totally-ordered, composing one of these queries with their supremum can amplify the gaps between them, leading to a violation of the commutativity condition. The proof of this technical result involves analyzing the topological structure of the hockey-stick curves of unordered queries. The details are given in Section˜8. Note that non-degeneracy is key as the following result fails for degenerate queries such as $L_{1}:=1/2\cdot 1_{-\ln{2}}+1/2\cdot 1_{\infty}$ and $L_{2}:=2/3\cdot 1_{0}+1/3\cdot 1_{\infty}$.
Lemma 4.
Let $L_{1}$ and $L_{2}$ be non-degenerate PLDs such that $L_{1}\npreceq L_{2}$ and $L_{2}\npreceq L_{1}$, i.e. $\{L_{1},L_{2}\}$ is not totally-ordered. Then, for either $i=1$ or $i=2$, we have
$L_{i}\oplus\sup\{L_{1},L_{2}\}\succ\sup\{L_{i}\oplus L_{1},L_{i}\oplus L_{2}\}.$
Proof of Theorem˜8.
The forward direction now follows directly from Lemma˜3 and Theorem˜7. As for the reverse direction, suppose that the subfamily of non-degenerate queries in $\mathcal{L}^{<k}$ is not totally-ordered, i.e. there exist $L_{1},L_{2}\in\mathcal{L}^{<k}$ that are not ordered. By Lemma˜4, we have that $L_{i}\oplus\sup\{L_{1},L_{2}\}\succ\sup\{L_{i}\oplus L_{1},L_{i}\oplus L_{2}\}$. On the other hand, if the commutativity condition of Theorem˜7 were satisfied for $\mathcal{L}^{<2k-3}$, then, letting $L_{i}=L_{i,1}\oplus\dots\oplus L_{i,t}$ ($t<k$), we would get
$L_{i}\oplus\sup\{L_{1},L_{2}\}$ $=L_{i,1}\oplus\dots\oplus L_{i,t}\oplus\sup\{L_{1},L_{2}\}$
$=L_{i,1}\oplus\dots\oplus L_{i,t-1}\oplus\sup\{L_{i,t}\oplus L_{1},L_{i,t}\oplus L_{2}\}$
$=\dots$
$=\sup\{L_{i,1}\oplus\dots\oplus L_{i,t}\oplus L_{1},L_{i,1}\oplus\dots\oplus L_{i,t}\oplus L_{2}\}$
$=\sup\{L_{i}\oplus L_{1},L_{i}\oplus L_{2}\},$
which is a contradiction. ∎
@@ 6 Notable Implications
6.1 Natural Filters Are NOT Free in General
In this section, we demonstrate examples of queries and budgets that do and do not pass the characterizing conditions of the previous section. For visualization purposes, it will be convenient to pass to tradeoff curves. Recall that analogous operation for the PLD supremum is the lower convex envelope (c.f. Proposition˜11). Recall also that the order of domination for PLDs is reversed for their tradeoff curves (see Proposition˜10). In this section, we will generally discuss piecewise linear tradeoff curves, which can be composed exactly as follows. A proof is given in Appendix˜B.
@@ 6.1 Natural Filters Are NOT Free in General
Consider two piecewise linear tradeoff curves $f_{1}$ and $f_{2}$ corresponding to PLDs $L_{1}$ and $L_{2}$. Let $(w_{i,j})_{j=1}^{n_{i}}$ be the lengths of the intervals over which $f_{i}$ is linear, let $(f^{\prime}_{i,j})_{j=1}^{n_{i}}$ be the slope of $f_{i}$ over those intervals, and let $\delta_{i}:=1-f_{i}(0)$. Now, let $(j_{k},\ell_{k})_{k=1}^{n_{1}n_{2}}$ be an enumeration of $n_{1}]\times[n_{2}]$ such that the product of the slopes $f^{\prime}_{1,j_{k}}\cdot f^{\prime}_{2,\ell_{k}}$ is descending in $k$, i.e. $f_{1,j_{1}}\cdot f_{2,\ell_{1}}$ is the product of the steepest slopes of $f_{1}$ and $f_{2}$. Then, the composed tradeoff curve $f$ corresponding to $L_{1}\oplus L_{2}$ is also a piecewise linear tradeoff curve over intervals of length $w_{k}:=w_{1,j_{k}}w_{2,\ell_{k}}$ with slopes $f^{\prime}_{k}:=-f^{\prime}_{1,j_{k}}\cdot f^{\prime}_{2,\ell_{k}}$ such that $1-f(0)=\delta_{1}+\delta_{2}-\delta_{1}\delta_{2}$.
An immediate implication of Theorem˜8 is that the natural filter is not universally free for general queries, even for a single adaptive round. This can be seen by noticing that $\mathcal{L}_{\mathrm{approx}}$ is not totally-ordered (recall Definition˜4). Figure˜2(a) demonstrates this visually by passing to the tradeoff curves of $\mathcal{L}_{\mathrm{approx}}$.
Corollary 1.
The natural filter for $\geq 2$ adaptive $(\varepsilon,\delta)$-DP queries is not universally free, i.e. there exists a budget $B\in\mathcal{L}_{\mathrm{all}}$ for which $\operatorname{PLD}(\operatorname{Nat}_{\mathcal{L}_{\mathrm{approx}},B}^{2})\npreceq B$.
(a) Queries $L_{1},L_{2}\in\mathcal{L}_{\mathrm{approx}}$ such that $\{L_{1},L_{2}\}$ is not totally-ordered and a budget $B\in\mathcal{L}_{\mathrm{approx}}$ such that $L_{1}\oplus L_{1},L_{1}\oplus L_{2}\preceq B$, yet $L_{1}\oplus\sup\{L_{1},L_{2}\}\npreceq B$.
(b) Queries $L_{1},L_{2}\in\mathcal{L}_{\mathrm{pure}}$ such that $\{L_{1},L_{2}\oplus L_{2}\}$ is not totally-ordered.
Figure 2: Counterexamples showing that natural filters fail for $(\varepsilon,\delta)$-DP queries and $(\varepsilon,\delta)$-DP budgets as well as $\varepsilon$-DP queries and general budgets.
Another interesting case is that of the $\varepsilon$-DP queries $\mathcal{L}_{\mathrm{pure}}$. Indeed, this family is totally-ordered and known to admit privacy filters under basic composition (Rogers et al., 2016). However, this family of queries is not closed under composition and $\mathcal{L}_{\mathrm{pure}}^{<3}$ is not totally-ordered anymore. Consequently, Theorem˜8 yields the following result, illustrated visually in Figure˜2(b).
Corollary 2.
The natural filter for queries $\mathcal{L}_{\mathrm{pure}}$ is universally free for $k=2$ queries but not $k\geq 3$ queries, i.e. $\operatorname{PLD}(\operatorname{Nat}_{\mathcal{L}_{\mathrm{pure}},B}^{2})\preceq B$ for all $B\in\mathcal{L}_{\mathrm{all}}$ whereas $\operatorname{PLD}(\operatorname{Nat}_{\mathcal{L}_{\mathrm{pure}},B}^{3})\npreceq B$ for some $B\in\mathcal{L}_{\mathrm{all}}$.
We have shown that natural filters are not universally free, even when restricted to $(\varepsilon,\delta)$-queries. These results are surprising, as all previous filter results were that filters are free (RDP/zCDP Feldman and Zrnic (2021); Lécuyer (2021), GDP Koskela et al. (2022); Smith and Thakurta (2022)) or asymptotically free ($(\varepsilon,\delta)$-DP Rogers et al. (2016); Whitehouse et al. (2023)). One might hope that restricting the budget would be sufficient to ensure free natural filters. Two natural candidates come to mind. First, restricting $B\in\mathcal{L}$—a practical condition of choosing privacy bounds from the legal query set itself. Second, we could restrict $B$ to a GDP bound $G_{\mu}$ or an approximate GDP bound $R_{0,\delta}\oplus G_{\mu}$ to support approximate $(\varepsilon,\delta)$-DP queries. These are good candidates, as the only known free natural filters Koskela et al. (2022); Smith and Thakurta (2022) are for $\mathcal{L}_{\mathrm{gdp}}$ queries and budgets. Unfortunately, neither restriction is sufficient.
The first point, that restricting the budget to $\mathcal{L}$ does not ensure the freeness of the natural filter, can be seen by applying Theorem˜6. Figure˜2(a) shows an example of queries $L_{1},L_{2}\in\mathcal{L}_{\mathrm{approx}}$ and a budget $B\in\mathcal{L}_{\mathrm{approx}}$ such that $L_{1}\oplus L_{1},L_{1}\oplus L_{2}\preceq B$ yet $L_{1}\oplus\sup\{L_{1},L_{2}\}\npreceq B$. In this case, Theorem˜6 implies $\operatorname{PLD}(\operatorname{Nat}_{\mathcal{L}_{\mathrm{approx}},B}^{2})\npreceq B$.
Corollary 3.
The natural filter is not free for $(\varepsilon,\delta)$-DP queries, even for a $(\varepsilon,\delta)$-DP budget, i.e. for some choice of $B\in\mathcal{L}_{\mathrm{approx}}$, we have that $\operatorname{PLD}(\operatorname{Nat}_{\mathcal{L}_{\mathrm{approx}},B}^{2})\npreceq B$.
(a) An example showing queries $L,L_{1},L_{2}\in\mathcal{L}_{\mathrm{all}}$ and a budget $G_{\mu}\in\mathcal{L}_{\mathrm{gdp}}$ such that $L\oplus L_{1},L\oplus L_{2}\preceq G_{\mu}$, yet $L\oplus\sup\{L_{1},L_{2}\}\npreceq G_{\mu}$.
(b) An example showing queries $L_{1},L_{2}\in\mathcal{L}_{\mathrm{approx}}$ and a budget $(\mu,\delta)$ such that $L_{1}\oplus L_{1},L_{1}\oplus L_{2}\preceq G_{\mu}\oplus R_{0,\delta}$, yet $L_{1}\oplus\sup\{L_{1},L_{2}\}\npreceq G_{\mu}\oplus R_{0,\delta}$.
Figure 3: Failure of natural filters for GDP and approximate GDP budgets.
Similarly, restricting the budget to a well-behaved class does not generally lead to free filters either. That GDP budgets do not suffice for general queries can again be shown by appealing to Theorem˜6 in the same manner, shown visually in Figure˜3(a). Likewise, approximate GDP budgets do not give free natural filters for $(\varepsilon,\delta)$-DP queries, as illustrated in Figure˜3(b).
Corollary 4.
The natural filter is not free for general queries, even for a GDP budget, i.e. for some $\mu\geq 0$, $\operatorname{PLD}(\operatorname{Nat}_{\mathcal{L}_{\mathrm{all}},G_{\mu}}^{2})\npreceq G_{\mu}$.
Furthermore, the natural filter is not free for $(\varepsilon,\delta)$-DP queries, even if the budget is chosen to be approximate GDP, i.e. for some $\mu\geq 0$ and $\delta\in[0,1]$, $\operatorname{PLD}(\operatorname{Nat}_{\mathcal{L}_{\mathrm{approx}},G_{\mu}\oplus R_{0,\delta}}^{2})\npreceq G_{\mu}\oplus R_{0,\delta}$.
6.2 Families of Queries With Universally Free Natural Filters
Theorem˜8 ensures that query families that are closed under composition and totally-ordered admit universally free natural filters.
In particular, since $G_{\mu_{1}}\oplus G_{\mu_{2}}=G_{\sqrt{\mu_{1}^{2}+\mu_{2}^{2}}}$ and since $G_{\mu_{1}}\preceq G_{\mu_{2}}$ whenever $0\leq\mu_{1}\leq\mu_{2}$, we have:
Corollary 5.
The natural filter for GDP queries is universally free, i.e. $\operatorname{PLD}(\operatorname{Nat}_{\mathcal{L}_{\mathrm{gdp}},B})\preceq B$ for all $B\in\mathcal{L}_{\mathrm{all}}$.
Likewise, as $R_{0,\delta_{1}}\oplus R_{0,\delta_{2}}=R_{0,\delta_{1}+\delta_{2}-\delta_{1}\delta_{2}}$ and as $R_{0,\delta_{1}}\preceq R_{0,\delta_{2}}$ for $0\leq\delta_{1}\leq\delta_{2}\leq 1$, we have:
Corollary 6.
The natural filter for $(0,\delta)$-DP queries is universally free, i.e. for every $B\in\mathcal{L}_{\mathrm{all}}$ we have $\operatorname{PLD}(\operatorname{Nat}_{\{R_{0,\delta:\delta\in[0,1]}\},B})\preceq B$.
7 Upper Bound for the Natural $(\varepsilon,\delta)$-DP PLD Filter
We consider the natural $(\varepsilon,\delta)$-DP PLD filter, which is simply an instantiation of Algorithm˜2 with the $(\varepsilon,\delta)$-DP budget $B=R_{\varepsilon,\delta}$ where, recalling Definition˜4,
$R_{\varepsilon,\delta}=\begin{cases}\infty&\text{w.p. }\delta\\ \varepsilon&\text{w.p. }\frac{1-\delta}{e^{-\varepsilon}+1}\\ -\varepsilon&\text{w.p. }\frac{1-\delta}{e^{\varepsilon}+1}\end{cases}.$
It is well known that a PLD $L$ is dominated by $R_{\varepsilon,\delta}$ iff $h_{L}(e^{\varepsilon})\leq\delta$. In other words, this filter uses the “natural” condition that checks whether $L_{1}\oplus\cdots\oplus L_{t}$ satisfies $(\varepsilon,\delta)$-DP.
Our main result is that, while the above natural $(\varepsilon,\delta)$-DP PLD filter can fail, it cannot fail too badly; in the sense that we still get a DP guarantee with a blow-up in parameters that is only polylogarithmic in $k,1/\delta$, as stated below [^2] Recall the notation $\operatorname{Nat}_{\mathcal{L},B}^{k}$ from Section 5.. We remark that our upper bound holds for the family $\mathcal{L}_{\mathrm{sym}}$ of all symmetric PLDs, which is a standard practical setting (see Appendix˜C).
Theorem 9.
Let $\varepsilon>0,\delta\in(0,1/4)$. Then, the natural $(\varepsilon,\delta)$-DP PLD filter for symmetric queries satisfies $\left({\chi\cdot(\varepsilon+1),\chi\cdot\delta}\right)$}-DP; in other words,
$\operatorname{PLD}{\left({\operatorname{Nat}_{\mathcal{L}_{\mathrm{sym}},R_{\varepsilon,\delta}}^{k}}\right)}\preceq R_{\chi\cdot(\varepsilon+1),\chi\cdot\delta},$
where
$\chi=O{\left({{\left({\log k+\log{\left({1+\frac{1}{\varepsilon}}\right)}}\right)}\cdot\log(1/\delta)}\right)}.$
When $\varepsilon=\Omega(1)$, the above guarantee is simply $(\chi^{\prime}\cdot\varepsilon,\chi^{\prime}\cdot\delta)$-DP where $\chi^{\prime}=O(\log k\cdot\log(1/\delta))$.
Remark 10.
In our current proof, the parameters are quite flexible and we can tradeoff the blow-up factors of the two parameters $\varepsilon,\delta$. For example, by changing $\varepsilon_{i}$ in Section˜7.1 to $\frac{2\varepsilon}{2^{n-1-i}}$., we can prove a guarantee of $\left(\chi\cdot\varepsilon,\chi\cdot\frac{\delta}{\varepsilon}\right)$-DP for $\chi=O(\log k\cdot\log(1/\delta))$. However, we choose to keep only a single setting of parameters in Theorem˜9 for simplicity.
For the same reason, we do not attempt to optimize $\chi$. In particular, it is plausible that our techniques can achieve $\chi=\tilde{O}(\log k+\log(1/\delta))$ with a more refined argument. Nevertheless, our proof cannot eliminate the dependency of $k$ or $\delta$ in $\chi$, and this remains an interesting open question.
Proof Overview.
We start with an informal overview of our proof, which consists of three steps:
  • Privacy Loss Discretization. Before we analyze the privacy filter, we start by providing a discretization argument for PLDs. Namely, we pick discretization points $\varepsilon_{0},\dots,\varepsilon_{n}$. Then, we “round” the PLD $L$ up or down to be these values (or their negations). We show that this process results in $L$ being “sandwiched” between two instances of “generalized randomized response”, which has the following form: Randomly select index $i$ from a certain probability distribution and then output $i$ together with the $\varepsilon_{i}$-DP randomized response. We show that $L$ is dominated by the “rounded up” (aka “pessimistic”) version of such randomized response while it dominates the “rounded down” (aka “optimistic”) version. We remark that our discretization argument is different than previous work (e.g. Meiser and Mohammadi (2018); Koskela et al. (2020); Doroshenko et al. (2022); Gopi et al. (2021); Ghazi et al. (2022)) which do not produce a generalized randomized response instance (especially for the pessimistic version). Indeed, we exploit the symmetry of $L$ to achieve our discretization.
  • Translation to Additive Filter. We consider the following “additive” filter, in which there is a non-negative real value budget associated with each $\varepsilon_{i}$ and each time we simply subtract from the budget; the filter rejects if we run out of budget for any $\varepsilon_{i}$. With the above discretization, we can show that the $(\varepsilon,\delta)$-PLD filter is “stronger” than such an additive filter (where the budget spent in each step is from the discretization). In other words, if one passes the PLD filter, then one also passes the additive filter. This means that, for the purpose of privacy guarantee, we may analyze the latter.
  • Analysis of Additive Filter via Post-Processing. In the final step, we provide the privacy analysis of the additive filter. Roughly speaking, we show that the adversary’s view of the additive filter can be shown to be a post-processing of a certain pre-specified number of randomized response (“with high probability”). Due to the post-processing property of DP, we can thus conclude that this is as private as composition of randomized response, which we then show to be $(\varepsilon^{\prime},\delta^{\prime})$-DP.
The above overview skips multiple subtle details, and the remainder of this section is devoted to carry out the above steps in full.
7.1 Discretization of PLD
Our discretization is based on a variant of randomized response where with a different prescribed probability, we run a randomized response with a different value of $\varepsilon$. We formalize this below.
Definition 11 (Randomized Response).
We use $\mathrm{RR}_{\varepsilon}$ to denote the $\varepsilon$-DP binary randomized response. That is, for $x\in\{0,1\}$,
$\mathrm{RR}_{\varepsilon}(x)=\begin{cases}x&\text{ with probability }\frac{1}{e^{-\varepsilon}+1},\\ 1-x&\text{ with probability }\frac{1}{e^{\varepsilon}+1}.\end{cases}$
Note that for $\varepsilon=\infty$, $\mathrm{RR}_{\varepsilon}(x)=x$ with probability one.
Let $\Delta_{n-1}:=\{\mathbf{p}=(p_{1},\dots,p_{n})\in\mathbb{R}_{\geq 0}\mid p_{1}+\cdots+p_{n}=1\}$ be the (n1)(n-1)-simplex, and let $\mathcal{E}_{n}:=\{\bm{\varepsilon}=(\varepsilon_{0},\dots,\varepsilon_{n})\in\mathbb{R}_{\geq 0}^{n+1}\mid 0=\varepsilon_{0}\leq\varepsilon_{1}\leq\cdots\leq\varepsilon_{n}=\infty\}$.
Definition 12 (Generalized Randomized Response).
For any $\mathbf{p}\in\Delta_{n-1}$ and $\bm{\varepsilon}\in\mathcal{E}_{n}$, we write $\mathrm{RR}_{\mathbf{p},\bm{\varepsilon}}$ to denote the following mechanism (where $x$ denotes the input):
  • Sample index $i\in[n]$ according to $\mathbf{p}$,
  • Sample $y\sim\mathrm{RR}_{\varepsilon_{i}}(x)$,
  • Output $(i,y)$.
We also define the following “rounded down” version $\mathrm{RR}^{\downarrow}_{\mathbf{p},\bm{\varepsilon}}$ of the above mechanism as follows where we use $\varepsilon_{0}$ to denote 0 throughout:
  • Sample index $i\in[n]$ according to $\mathbf{p}$,
  • Sample $y\sim\mathrm{RR}_{\varepsilon_{i-1}}(x)$,
  • Output $(i,y)$.
Note that $\mathrm{RR}_{\mathbf{p},\bm{\varepsilon}}$ and $\mathrm{RR}^{\downarrow}_{\mathbf{p},\bm{\varepsilon}}$ only differ on the second step, where the former invokes $\mathrm{RR}_{\varepsilon_{i}}$ but the latter invokes $\mathrm{RR}_{\varepsilon_{i-1}}$.
In addition to Generalized Randomized Response, another notion we will use for the discretization step is the notion of discretized privacy loss profile, which is a vector $\mathbf{p}=(p_{1},\dots,p_{n})$ where $p_{i}$ is the probability that the privacy loss falls into $\varepsilon_{i-1},\varepsilon_{i})$ or $(-\varepsilon_{i},\varepsilon_{i}]$. This is formalized below.
Definition 13 (Discretized Privacy Loss Profile).
For any PLD $L$, the discretized privacy loss profile of $L$ and $\bm{\varepsilon}\in\mathcal{E}_{n}$, denoted by $\mathrm{disc}_{\bm{\varepsilon}}(L)$, is the vector $\mathbf{p}=(p_{1},\dots,p_{n})\in\Delta_{n-1}$ where
$p_{i}=\begin{cases}\Pr_{Z\sim L}\left[\varepsilon_{i-1}\leq|Z|<\varepsilon_{i}\right]&\text{ if }i\neq n,\\ \Pr_{Z\sim L}\left[\varepsilon_{n-1}\leq|Z|\right]&\text{ if }i=n.\end{cases}$
The “optimistic” and “pessimistic” discretization of $L$ is exactly based on $\mathrm{RR}^{\downarrow}_{\mathbf{p},\bm{\varepsilon}}$ and $\mathrm{RR}_{\mathbf{p},\bm{\varepsilon}}$, and we have the following theorems which gives the aforementioned “sandwich” around the pair $L$.
Theorem 11 (Pessimistic Estimate).
Suppose that $L$ is a symmetric PLD, and let $\mathbf{p}$ be its discretized privacy loss profile. Then, $\operatorname{PLD}(\mathrm{RR}_{\mathbf{p},\bm{\varepsilon}}(0),\mathrm{RR}_{\mathbf{p},\bm{\varepsilon}}(1))\succeq L$.
Theorem 12 (Optimistic Estimate).
Suppose that $L$ is a symmetric PLD, and let $\mathbf{p}$ be its discretized privacy loss profile. Then, $\operatorname{PLD}(\mathrm{RR}^{\downarrow}_{\mathbf{p},\bm{\varepsilon}}(0),\mathrm{RR}^{\downarrow}_{\mathbf{p},\bm{\varepsilon}}(1))\preceq L$.
The remainder of this subsection are devoted to the proofs of these two theorems.
7.1.1 Useful Lemmata
Before we can prove Theorems˜11 and 12, we will first prove a couple of technical lemmata that would be useful. First is the symmetry of $\mathrm{RR}$ and $\mathrm{RR}^{\downarrow}$, which are simple to prove.
Lemma 5.
For any $\mathbf{p},\bm{\varepsilon}$, the pairs $(\mathrm{RR}_{\mathbf{p},\bm{\varepsilon}}(0),\mathrm{RR}_{\mathbf{p},\bm{\varepsilon}}(1))$ and $(\mathrm{RR}^{\downarrow}_{\mathbf{p},\bm{\varepsilon}}(0),\mathrm{RR}^{\downarrow}_{\mathbf{p},\bm{\varepsilon}}(1))$ are symmetric.